in triangle abc side ab is 2 cm shorter than side ac, while side bc is 1cm longer than ac. if the perimeter is 210 cm. how long are sides? show linear equation and solution

Answer by addingup(3677)

abc, we have three sides:
ab,ac, bc
-----------------
ab = ac-2
bc = ac+1
The perimeter is:
ab+ac+bc = 210 now substitute for ab and bc:
ac-2+ac+1+ac = 210
3ac-1 = 210
3ac = 211
ac = 70.333
------------------------
Check:
(70.333-2)+(70.333+1)+70.333 = 210 Correct